OpenStudy (anonymous):

Help please!! ((: Solve each equation by factoring:

OpenStudy (anonymous):

\[6n^2-18n-18=6\]

OpenStudy (mathstudent55):

First, subtract the 6 from the right side to teh left side.

OpenStudy (mathstudent55):

Then divide the entire equation by 6.

OpenStudy (anonymous):

so: \[6n^2-18n-24=0\]

OpenStudy (anonymous):

or would it be positive 24 b/c the 2 negatives?

jimthompson5910 (jim_thompson5910):

-18 - 6 = -24 so you have it correct

OpenStudy (mathstudent55):

You are correct. Now, since 6, -18, and -24 have a common facotr of 6, divide both sides by 6.

OpenStudy (anonymous):

Mmkay thanx!

OpenStudy (mathstudent55):

Did you finish?

OpenStudy (mathstudent55):

6n^2 - 18n - 24 = 0 Now we divide both sides by 6 since 6 is the GCF of all terms: n^2 - 3n - 4 = 0 Now we factor the left side: (n - 4)(n + 1) = 0 Set each factor equal to zero: n - 4 = 0 or n + 1 = 0 n = 4 or n = -1

OpenStudy (anonymous):

Thanx!!! Sorry, I was busy with something else! Thanx! ((:

OpenStudy (mathstudent55):

wlcm